OpenStudy (anonymous):

Which statement is true about the product square root of 2(3square root of 2 + square root of 14)? It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

geerky42 (geerky42):

\[\sqrt{2}(3\sqrt2 + \sqrt{14})\]Right? What do you think so far?

geerky42 (geerky42):

You can split \(\sqrt{14}\) into two square roots; \(\sqrt{14} = \sqrt{2\cdot7}=\sqrt2\sqrt7\)

geerky42 (geerky42):

\[\sqrt{2}(3\sqrt2 + \sqrt2\sqrt7)\] Now distribute \(\sqrt2\)

OpenStudy (anonymous):

Um I think I get it but what next?

OpenStudy (anonymous):

simplifying further, you'll get \[6+2\sqrt{7}\]

OpenStudy (anonymous):

so would that be c ? @aryandecoolest

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

C